COLLISIONAL ORIENTATION TRANSFER FACILIATED POLAROZATION SPECTROSCOPY^{a}

\underline{JIANMEI BAI}, E.H.AHMED, B. BESER, Y. GUAN, A. M. LYYRA, Temple University; S. ASH-MAN, C. M. WOLFE, J. HUENNEKENS, Lehigh University.

Collisional orientation transfer facilitated V-type double-resonance polarization spectroscopy technique was applied to study the A-b complex of Rb_{2}^{a} and Cs_{2}^{b}. Since spectral congestion makes it difficult to find isolated pump transitions for heavy molecules such as Rb_{2} and Cs_{2}, this technique significantly enlarges the range of rotational levels that can be observed per vibrational level. Collisional satellite lines with ΔJ up to 58 were observed in the Rb_{2} polarization experiment. In the Cs_{2} experiment, due to weaker Franck-Condon factors, collisional satellite lines with ΔJ_{max} equals to 12 were observed. Collisional orientation transfer in polarization spectroscopy was first observed with buffer gas pressure of several hundred Torr^{c}. The high pressure led to loss of spectral resolution from collisional broadening. Only 1 to 3 Torr of argon buffer gas pressure was used in our experiments to obtain spectra with much higher resolution. Among the six types of possible probe signals^{d}, we assigned and analyzed the signals from the V type excitation scheme. The data was used in the global deperturbation analysis of the A-b complex of both Rb_{2} and Cs_{2}.

aFunded by NSF PHY 0555608 and PHY 0855502

aH. Salami et al. Phys. Rev. A 80, 022515 (2009)

bJianmei Bai et al., Phys. Rev. A, to appear (2011)

cB. Teets et al. Phys. Rev. Lett. 37, 683 (1976)

dN. Okada et al. J. Chem. Phys. 105, 3458 (1996)